OpenStudy (anonymous):

solve using addition method 2x-y=8 and 4x+3y=36

OpenStudy (anonymous):

plzzzz help natala

OpenStudy (anonymous):

frist eqution time 3 u get 6x -3y =24 frist eqution 4x +3y =36 second keep the same

OpenStudy (anonymous):

ty will u keep helping me with my problems

OpenStudy (anonymous):

TY for the medal

OpenStudy (anonymous):

wc[;

OpenStudy (anonymous):

god cluck with your home work

OpenStudy (anonymous):

waite

OpenStudy (anonymous):

I caculer wrong

OpenStudy (anonymous):

with wat part

OpenStudy (anonymous):

6x -3y =24 Add 4x +3y =36 -------------------- 10x = 60 x = 66/10 x =6

OpenStudy (anonymous):

u put that already

OpenStudy (anonymous):

subtituion x= 6 solve for y (4*6)+3y=36 24 + 3y =36 3y =36-24 y = 12\3 y = 4

OpenStudy (gw2011):

x=6 and y=4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh i get wat u did its cool we all make mistakes

OpenStudy (anonymous):

now we have corect answer

OpenStudy (anonymous):

ty im typing up new / will u help me with it

OpenStudy (anonymous):

I try

OpenStudy (anonymous):

k just a sec

OpenStudy (anonymous):

k its posted on the ? page

OpenStudy (anonymous):

ok